OpenStudy (anonymous):

In a first order process, it was found that 1/7 of a sample had been consumed in 20 minutes. what is the half life of this reaction in minutes?

OpenStudy (aaronq):

you can use the first order integrated rate law: \(ln[A]_t=-kt+ln[A]_o\) to find the rate constant, k, then use: \(t_{1/2}=\dfrac{ln(2)}{k}\) to find the half-life.